# arXiv:2403.09539v2[cs.CL] 15 Mar 2024

### v2

@@ Abstract
The commercialization of large language models (LLMs) has led to the common practice of restricting access to proprietary models via an limited API. In this work we show that, with a conservative assumption about the model architecture, it is possible to learn a surprisingly large amount of non-public information about an API-protected LLM from a small number of API queries (e.g., costing under \$1,000 for OpenAI’s gpt-3.5-turbo). Our findings are centered on one key observation: most modern LLMs suffer from a softmax bottleneck, which restricts the model outputs to a linear subspace of the full output space. We exploit this fact to unlock several capabilities: efficiently discovering the LLM’s hidden size, obtaining cheap full-vocabulary outputs, detecting and disambiguating different model updates, identifying the source LLM given a single full LLM output, and even estimating the output layer parameters. Our empirical investigations show the effectiveness of our methods, which allow us to estimate the embedding size of OpenAI’s gpt-3.5-turbo to be about 4,096. Lastly, we discuss ways that LLM providers can guard against these attacks, as well as how these capabilities can be viewed as a feature (rather than a bug) by allowing for greater transparency and accountability.
@@ 1 Introduction
As large language models (LLMs) have become more capable and valuable, it has become increasingly common for companies and organizations to train closed-source LLMs and make them available only via an API (e.g., OpenAI et al., 2024). This setup may foster a false sense of security for LLM providers, who might mistakenly assume that information about their model architecture is private, and that certain types of attacks on their LLM are infeasible. On the flip side, users must seemingly take providers’ word that LLMs only change when the provider publicly announces version updates.
@@
Figure 1: LLM outputs are constrained to a low-dimensional subspace of the full output space. We can use this fact to glean information about API-protected LLMs by analyzing their outputs. Here we show how a toy LLM’s low-dimensional embeddings in $\mathbb{R}^{d}$ (illustrated here as a 1-D space) are transformed linearly into logits in $\mathbb{R}^{d}$ (here, a 3D space) via the softmax matrix $\boldsymbol{W}$. The resulting outputs lie within a ($d=1$)-dimensional subspace of the output space. The softmax function also preserves this low-dimensionality when transforming the logits into probabilities distributions on $\Delta_{v}$. We call this low-dimensional subspace the image of the model. We can obtain a basis for the image of an API-protected LLM by collecting $d$ of its outputs. The LLM’s image can reveal non-public information, such as the LLM’s embedding size, but it can also be used for accountability, such as verifying which LLM an API is serving.
Despite this apparent state of affairs, in reality many LLM APIs reveal much more information about their underlying models than previously thought. In this paper we show how to extract detailed information about LLM parameterization using only common API configurations. Our findings allow LLM clients to hold providers accountable by tracking model changes, recovering hidden prompts and cheaply obtaining full vocabulary outputs. At the same time, our approach also allows LLM providers to establish unique identities for their models, enabling trust with their users as well as improved accountability. Table 1 summarizes the applications that we explore in this paper.
Table 1: An overview of our proposed applications that exploit LLM images.
Application Section
Efficiently extract full LLM outputs. 3
Find the embedding size of an LLM and guess its parameter count. 4
Identify which LLM produced a given output. 5
Detect when and what type of LLM updates occur. 5.1
Find tokenization bugs (unargmaxable tokens). 6.1
Approximately reconstruct the LLM’s softmax matrix. 6.2
Cheaply and accurately reconstruct “hidden prompts”. 6.3
Implement basis-aware decoding algorithms. 6.4
Our method exploits the low-rank output layer common to most LLM architectures by observing that this layer’s outputs are constrained to a low-dimensional subspace of the full output space, as illustrated in Figure 1. We call this restricted output space the LLM’s image. We can obtain a basis for this image by collecting a small number of LLM outputs, and we develop novel algorithms that enable us to accomplish this cheaply and quickly for standard LLM APIs.
Obtaining the LLM image (§2) allows to explore several capabilities that cover a broad set applications, and empirically verify several of them. For instance, we design and empirically verify an algorithm for finding the embedding size of an API-based LLM, and demonstrate the effectiveness of our method by estimating the embedding size of gpt-3.5-turbo, a closed-source API-protected LLM, to be 4,096. We also empirically show effectiveness of using LLM images as unique signatures that can be used to identify outputs from a model with high accuracy, a useful property for API LLM accountability. The sensitivity of these signatures to slight changes in the LLM parameters also makes them suitable for inferring granular information about model parameter updates.
@@
Considering several proposals to mitigate this vulnerability, we find no obvious fix to prevent obtaining LLM images without dramatically altering the LLM architecture. While providers may choose to alter the API to hide this information, the relevant API features have valuable and safe use cases for the LLM clients, who may rely on access to features like logit bias. Though our findings be viewed as a warning to LLM providers to carefully consider the consequences of their LLM architectures and APIs, we prefer to view our findings as a potential feature that LLM providers may choose to keep in order to better maintain trust with their customers by allowing outside observers to audit their model.
@@ 2 LLM outputs are restricted to a low-dimensional linear space
Consider the architecture of a typical LLM, depicted in Figure 2. A Transformer with embedding size $d$ outputs a low-dimensional contextualized embedding $\boldsymbol{h}\in\mathbb{R}^{d}$ (or simply embedding). Projecting the embedding onto $\mathbb{R}^{v}$ via the linear map defined by the LLM’s softmax matrix $\boldsymbol{W}$, we obtain the logits $\boldsymbol{\ell}=\boldsymbol{W}\boldsymbol{h}$. Because $\boldsymbol{W}$ is in $\mathbb{R}^{v\times d}$, its rank (i.e., number of linearly independent columns) is at most $d$. The rank of a matrix corresponds to the dimension of the image of the linear map it defines, i.e., the vector space comprising the set of possible outputs of the function. In other words, if the linear map $w$ is defined as $w(\boldsymbol{h})=\boldsymbol{W}\boldsymbol{h}$, then $w$’s image $\operatorname{im}(w)=\{w(\boldsymbol{h})\in\mathbb{R}^{v}:\boldsymbol{h}\in\mathbb{R}^{d}\}$ is a $d$-dimensional subspace of $\mathbb{R}^{v}$. Thus, the LLM’s logits will always lie on the $d$-dimensional [^1] More accurately, the logits will always lie on an at-most-$d$-dimensional subspace. For convenience, we assume full-rank matrices, and thus a $d$-dimensional subspace. subspace of $\mathbb{R}^{v}$.
Refer to caption Refer to caption Refer to caption Refer to caption
@@
We now turn our attention to the LLM’s final output: the next-token distribution $\boldsymbol{p}$. Due to the softmax function, this is a valid probability distribution over $v$ items, meaning it is a $v$-tuple of real numbers between 0 and $1$ whose sum is $1$. The set of valid probability distributions over $v$ items is referred to as the $v$-simplex, or $\Delta_{v}$. Perhaps surprisingly, $\Delta_{v}$ is also a valid vector space (albeit under a non-standard definition of addition and scalar multiplication) and is isomorphic to a (v1)(v-1)-dimensional vector subspace of $\mathbb{R}^{v}$ (Aitchison, 1982; Leinster, 2016). In particular, it is isomorphic to the hyperplane $U_{v}$ that is perpendicular to the all-ones vector $\mathbf{1}_{v}$, as illustrated in Figure 3. The softmax function is thus an isomorphism $U_{v}\to\Delta_{v}$ whose inverse mapping $\Delta_{v}\to U_{v}$ is the center log ratio transform (clr), which is defined as
$\operatorname{clr}(\boldsymbol{p})=\log\frac{\boldsymbol{p}}{\left(\prod_{i=1}^{v}p_{i}\right)^{\frac{1}{v}}}\,=\log\boldsymbol{p}-\frac{1}{v}\sum_{i=1}^{v}\log p_{i}.$ (1)

### v3

# arXiv:2403.09539v3[cs.CL] 08 Nov 2024

@@ Abstract
Large language model (LLM) providers often hide the architectural details and parameters of their proprietary models by restricting public access to a limited API. In this work we show that, with only a conservative assumption about the model architecture, it is possible to learn a surprisingly large amount of non-public information about an API-protected LLM from a relatively small number of API queries (e.g., costing under \$$1000$ USD for OpenAI’s gpt-3.5-turbo). Our findings are centered on one key observation: most modern LLMs suffer from a softmax bottleneck, which restricts the model outputs to a linear subspace of the full output space. We exploit this fact to unlock several capabilities, including (but not limited to) obtaining cheap full-vocabulary outputs, auditing for specific types of model updates, identifying the source LLM given a single full LLM output, and even efficiently discovering the LLM’s hidden size. Our empirical investigations show the effectiveness of our methods, which allow us to estimate the embedding size of OpenAI’s gpt-3.5-turbo to be about $4096$. Lastly, we discuss ways that LLM providers can guard against these attacks, as well as how these capabilities can be viewed as a feature (rather than a bug) by allowing for greater transparency and accountability.
@@ 1 Introduction
As large language models (LLMs) become more capable and valuable, many LLM providers have shifted toward training closed-source production LLMs, accessible only via (paywall-restricted) APIs to protect trade secrets (e.g., OpenAI et al., 2024). As a useful feature, these APIs often allow users to access the probabilities (or logits, i.e., unnormalized scores) that the LLM assigns to specific tokens. It turns out that such an interface reveals much more information about the underlying model than previously thought, including exposing non-public architectural details of the LLM. Closed APIs, therefore, may provide a false sense of security to production LLM providers who may assume that their product information is private. At the same time, these capabilities empower the community with tools to audit LLM providers for bad behaviors such as unannounced model updates and abuse of open-source LLMs (Mökander et al., 2023).
@@
Figure 1: LLM outputs are constrained to a low-dimensional subspace of the full output space. We can use this fact to glean information about API-protected LLMs by analyzing their outputs. Here we show how a toy LLM’s low-dimensional embeddings in $\mathbb{R}^{d}$ (illustrated here as a 1-D space) are transformed linearly into logits in $\mathbb{R}^{v}$ (here, a 3D space) via the softmax matrix $\boldsymbol{W}$. The resulting outputs lie within a ($d=1$)-dimensional subspace of the output space. We call this low-dimensional subspace the image of the model. We can obtain a basis for the image of an API-protected LLM by collecting $d$ of its outputs. The LLM’s image can reveal non-public information, such as the LLM’s embedding size, but it can also be used for accountability, such as verifying which LLM an API is serving.
In this paper we show that is possible to extract detailed information about LLM parameterization using only common API configurations. Our findings are centered on one key observation: most modern LLMs suffer from a softmax bottleneck (Yang et al., 2018), which restricts the model outputs to a linear subspace of the full output space, as illustrated in figure˜1. We call this restricted output space the LLM’s image2). This image can be obtained by collecting a small number of LLM outputs, and serves as a unique identifier or a signature for the model. We propose several algorithms that enable us to obtain the LLM image at low cost and speed for standard LLM APIs (§3).
Furthermore, we show that LLM images are useful in several applications, revealing important information about the model architecture. Concretely, we can exploit this image to provide an efficient algorithm for extracting full LLM outputs (§4), to find the hidden dimension (embedding) size of the LLM (§5), to detect and identify different kind of updates made to the model and to attribute model outputs to a specific model (§6), among other potential applications (§7). We demonstrate several of our proposed applications in order to empirically verify their effectiveness. Notably, we design and implement an algorithm for finding the embedding size of an API-based LLM, and use it to estimate the embedding size of gpt-3.5-turbo (a closed-source API-protected LLM) to be $4096$.
Overall, our proposed methods have benefits for both LLM providers and for their clients. LLM providers may use their model images to establish unique identities for their models, thereby protecting their product and building trust with clients. Our proposed methods will also allow LLM clients to hold providers accountable for malicious behavior (Anderljung et al., 2023). As a concrete use case in accountability, we demonstrate, using checkpoints from open-source LLMs that our LLM images can be used to attribute LLM output probabilities (or logits) to their generating model with high accuracy. The sensitivity of our LLM image to slight changes in the LLM parameters also makes them suitable for inferring granular information about the specific type of model update.
@@
Considering several proposals to guard access to LLM images, we find no obvious fix without dramatically altering the LLM architecture or making the API considerably less useful (§8). Providers who choose to alter their API to prevent LLM image access risk removing interfaces with valuable and safe use cases for LLM clients. Though our findings could be viewed as a bug that LLM providers might feel compelled to patch, we prefer to view them as features that LLM providers may choose to keep in order to better maintain trust with their customers by allowing outside observers to audit their model. Ultimately, our results serve as a recommendation to LLM providers to carefully consider the consequences of their LLM architectures and APIs.
This paper’s contributions include
  • A method for extracting information about API-protected models, including the model’s output space and embedding size.
  • Methods for extracting full-vocabulary logprob outputs from LLM APIs.
  • An estimate of the embedding size of an API-protected LLM (gpt-3.5-turbo).
  • An accelerated logprob extraction algorithm based on the LLM image.
  • An exploration of several other applications of our method for model accountability.
Concurrently with our work, Carlini et al. (2024) propose a very similar approach for exposing details of production LLMs, though with a focus on defenses and mitigations against such attacks. The “final layer” that they extract in their attack corresponds to what we refer to in our paper as the model image. We view our papers as complementary, since our work emphasises practical applications of LLM images for better LLM accountability.
@@ 2 LLM outputs are restricted to a low-dimensional linear space
The outputs of typical LLMs are naturally constrained to lie on a $d$-dimensional subspace of the full output space (Yang et al., 2018; Finlayson et al., 2024). To understand this, begin by considering the architecture of a typical LLM (figure˜2). In this architecture, a transformer [^1] Technically, any neural network suffices here. with embedding size $d$ outputs a low-dimensional contextualized embedding $\boldsymbol{h}\in\mathbb{R}^{d}$ (or simply embedding). Projecting the embedding onto $\mathbb{R}^{v}$ via the linear map defined by the LLM’s softmax matrix $\boldsymbol{W}$, we obtain logits $\boldsymbol{\ell}=\boldsymbol{W}\boldsymbol{h}$.
Theorem 1 (Low-rank logits).
LLM logits lie on a $d$-dimensional subspace of $\mathbb{R}^{v}$.
Proof.
Because $\boldsymbol{W}$ is in $\mathbb{R}^{v\times d}$, its rank (i.e., number of linearly independent columns) is at most $d$. The rank of a matrix corresponds to the dimension of the image of the linear map it defines, i.e., the vector space comprising the set of possible outputs of the function. In other words, if the linear map $w$ is defined as $w(\boldsymbol{h})=\boldsymbol{W}\boldsymbol{h}$, then $w$’s image $\operatorname{im}(w)=\{w(\boldsymbol{h})\in\mathbb{R}^{v}:\boldsymbol{h}\in\mathbb{R}^{d}\}$ is a $d$-dimensional subspace of $\mathbb{R}^{v}$. ∎
Thus, the LLM’s logits will always lie on the $d$-dimensional [^2] More accurately, the logits will always lie on an at-most-$d$-dimensional subspace. For convenience, we assume full-rank matrices, and thus a $d$-dimensional subspace. subspace of $\mathbb{R}^{v}$.
@@
The low-dimensional restriction of logits actually translates into a similar restriction on LLM probabilities, which is useful since most LLM APIs return (log-)probabilities instead of logits. To understand why this is, we turn our attention to the model’s next-token distribution $\boldsymbol{p}=\operatorname{softmax}(\boldsymbol{\ell})$. The softmax function [^3] Our use of “softmax” here corresponds to temperature softmax with temperature 1. transforms logits into a valid probability distribution, i.e., a $v$-tuple of real numbers between 0 and $1$ whose sum is $1$. The set of valid probability distributions over $v$ items is commonly referred to as the $v$-simplex, or $\Delta_{v}$.
Theorem 2 (Low-rank probabilities).
LLM probabilities lie on a $d$-dimensional subspace of $\Delta_{v}$.
Proof.
Perhaps surprisingly, $\Delta_{v}$ is also a valid (v1)(v-1)-dimensional vector space (albeit under non-standard definitions of addition and scalar multiplication) and the softmax function is a linear map $\mathbb{R}^{v}\to\Delta_{v}$ (Aitchison, 1982; Leinster, 2016). Therefore, since $\dim(\operatorname{im}(w))=d$, we also know that $\operatorname{im}(\operatorname{softmax}\circ w)$ is a $d$-dimensional subspace of $\Delta_{v}$. ∎
Thus, LLM output probabilities must also reside in a $d$-dimensional subspace.
Finally, since $\Delta_{v}$ is an unconventional vector space, it is useful to translate LLM probability distributions into a more conventional vector space. To do so, we use the fact that $\Delta_{v}$ is isomorphic to a vector subspace of $\mathbb{R}^{v}$ via the softmax function. In particular, figure˜3 illustrates how $\Delta_{v}$ is isomorphic to the hyperplane $U_{v}$ that is perpendicular to the all-ones vector $\mathbf{1}_{v}$. The isomorphism’s inverse mapping $\Delta_{v}\to U_{v}$ is the center log ratio transform $\operatorname{clr}(\boldsymbol{p})=\log\boldsymbol{p}-\frac{1}{v}\sum_{i=1}^{v}\log p_{i}.$ By linearity, $\operatorname{im}(\operatorname{clr}\circ\operatorname{softmax}\circ w)$ is a $d$-dimensional subspace of $U_{v}\subset\mathbb{R}^{v}$.
@@
Thus, LLM outputs occupy $d$-dimensional subspaces of logit space $\mathbb{R}^{v}$, probability space $\Delta_{v}$, and $U_{v}$. We call these subspaces the image of the LLM on each given space. A natural consequence this low-dimensionality is that any collection of $d$ linearly independent LLM outputs form a basis for the image of the model, i.e., all LLM outputs can be expressed as a linear combination of these outputs.
3 Obtaining full outputs from API-protected LLMs
As explained in the previous section, in order to obtain the image of an LLM, we need to collect outputs (token probabilites for each token in the vocabulary) from the LLM. Unfortunately, most LLM APIs do not return full outputs, likely because full outputs are large and expensive to send over an API, but perhaps also to prevent API abuse, since full LLM outputs contain lots of useful information (Dosovitskiy and Brox, 2016; Morris et al., 2023) and can be used to distill models (e.g., Hinton et al., 2015; Hsieh et al., 2023). In their paper, Morris et al. (2023) give an algorithm for recovering full outputs from restricted APIs by taking advantage of a common API option that allows users to add a bias term $\beta\leq\beta_{\max}$ to the logits for specific tokens. The algorithm they describe requires $v\log(\beta/\epsilon)$ calls to the API to obtain one full output with precision $\epsilon$, or exactly $v$ calls when the API returns log-probabilities of the top-2 tokens. Carlini et al. (2024) also propose a logprob-free method based on the same principle.
We give a variant of their algorithm for APIs that return the log-probability of the top-$k$ tokens that obtains full outputs in $v/k$ API calls. We find that this improved algorithm suffers from numerical instability, and give a numerically stable algorithm that obtains full outputs in v/(k1)v/(k-1) API calls. We also give a practical algorithm for dealing with stochastic APIs that randomly choose outputs from a set of $n$ possible outputs, a behavior we observe in OpenAI’s API. This algorithm allows the collection of full outputs in nv/(k2)nv/(k-2) API calls on average. table˜1 gives an overview of our algorithms with back-of-the envelope cost estimates for a specific LLM.
Algorithm Complexity API calls per output Image price (USD)
Logprob-free (Morris et al., 2023) $v\log(\beta_{\max}/\epsilon)$ 800 000800\,000 16 38416\,384
With logprobs $v/k$ 20 00020\,000 $410$
Numerically stable v/(k1)v/(k-1) 25 00025\,000 $512$
Stochastically robust nv/(k2)nv/(k-2) 133 000133\,000 $2724$
LLM Image (§4) $O(d)$ $800$
Table 1: A summary of our proposed algorithms for obtaining full LLM outputs, with estimates for the number of API calls required per output, and the price of acquiring the model image. Estimates are based on a gpt-3.5-turbo-like API LLM with v=100 000v=\$100\,000\$, $d=4096$, $\epsilon=10^{-6}$, $k=5$, $\beta_{\max}=100$, and $n=4$. Note that the $O(d)$ algorithm cannot be used to obtain the LLM image, since it relies on having LLM image as a preprocessing step.
3.1 Full outputs from APIs with logprobs
Our goal is to recover a full-vocabulary next-token distribution $\boldsymbol{p}\in\Delta_{v}$ from an API-protected LLM. We will assume that the API accepts a prompt on which to condition the distribution, as well as a list of up to $k$ tokens and a bias term $\beta\leq\beta_{\max}$ to add to the logits of the listed tokens before applying the softmax function. The API returns a record with the $k$ most likely tokens and their probabilities from the biased distribution. For instance, querying the API with $k$ maximally biased tokens, which (without loss of generality) we will identify as tokens $1,2,\ldots,k$, yields the top-$k$ most probable tokens from the biased distribution $\boldsymbol{p}^{\prime}=\operatorname{softmax}(\boldsymbol{\ell}^{\prime})$ where
$\ell^{\prime}_{i}=\begin{cases}\ell_{i}+\beta_{\max}&i\in\{1,2,\ldots,k\}\\ \ell_{i}&\text{otherwise}\end{cases}$ (1)
Assuming that the logit difference between any two tokens is never greater than $\beta_{\max}$, these top-$k$ biased probabilities will be $p^{\prime}_{1},p^{\prime}_{2},\ldots,p^{\prime}_{k}$. For each of these biased probabilities $p^{\prime}_{i}$, we can solve for the unbiased probability as
$p_{i}=\frac{p^{\prime}_{i}}{\exp\beta_{\max}-\exp\beta_{\max}\sum_{j=1}^{k}p^{\prime}_{j}+\sum_{j=1}^{k}p^{\prime}_{j}}$ (2)
(proof in the §A.1). Thus, for each API call, we can obtain the unbiased probability of $k$ tokens, and obtain the full distribution in $v/k$ API calls.
3.2 Numerically stable full outputs from APIs
In practice, the algorithm described in §3.1 suffers from severe numerical instability, which can be attributed to the fast-growing exponential term $\exp\beta_{\max}$, and the term $\sum_{j=1}^{k}p^{\prime}_{j}$ which quickly approaches 1. We can eliminate the instability by sacrificing some speed and using a different strategy to solve for the unbiased probabilities. Without loss of generality, let $p_{v}$ be the maximum unbiased token probability. This can be obtained by querying the API once with no bias. If we then query the API and apply maximum bias only to tokens $1,2,\ldots,1-k$, then the API will yield $p^{\prime}_{1},p^{\prime}_{2},\ldots,p^{\prime}_{k-1}$ and $p^{\prime}_{v}$. We can then solve for the unbiased probabilities of the k1k-1 tokens
$p_{i}=\exp(\log p^{\prime}_{i}-\beta_{\max}-\log p^{\prime}_{v}+\log p_{v})$ (3)
(proof in §A.2). By finding k1k-1 unbiased token probabilities with every API call, we obtain the full output in v/(k1)v/(k-1) calls total. Algorithm 1 gives a formal description of this method.
Algorithm 1 Our numerically stable probability extraction algorithm, which takes a set of contexts, a set of tokens, and an API for an LLM with vocabulary $\mathcal{V}$ and returns the LLM’s probabilities for each token in each context. The API takes a context and a set of token-bias pairs and returns the top-$k$ biased probabilities. $\mathcal{P}(S)$ denotes the power set of a set $S$.
function Extract( $\mathsf{contexts}\subseteq\mathcal{V}^{*}$, $\mathsf{tokens}\subseteq\mathcal{V}$, $\mathsf{API}:\mathcal{V}^{*}\times\mathcal{P}(\mathcal{V}\times\mathbb{R})\to\mathcal{P}(\mathcal{V}\times\mathbb{R})$ )
  $\mathsf{probs}\in\mathbb{R}^{\lvert\mathsf{contexts}\rvert\times\lvert\mathsf{tokens}\rvert} \triangleright$ Initialize empty index to collect probabilities
  for $c\in\mathsf{contexts}$ do
   $(v,p_{v})=\operatorname{argmax}_{(i,p)\in\mathsf{API}(c,\emptyset)}p$ $\triangleright$ Get top probability token
   for each batch $T\subseteq\mathsf{tokens}$ with $\lvert T\rvert=k-1$ do $\triangleright$ Partition $\mathsf{tokens}$ into batches
     $\mathsf{bias}=\{(i,\beta)\}_{i\in T}$ $\triangleright$ Set biases to $\beta$ for batch tokens
     $\{(v,p^{\prime}_{v})\}\cup\{(i,p^{\prime}_{i})\}_{i\in T}=\mathsf{API}(c,\mathsf{bias})$ $\triangleright$ Get biased probabilities
     for $i\in T$ do
      $\mathsf{probs}_{c,i}=\exp(\log p^{\prime}_{i}-\beta-\log p^{\prime}_{v}+\log p_{v})$ $\triangleright$ Get unbiased probability
     end for
   end for
  end for
  return $\mathsf{probs}$
end function
3.3 Full outputs from stochastic APIs
Each of the above algorithms assume that the API is deterministic, i.e., the same query will always return the same output. However, this may not always be the case; for instance, we find that OpenAI’s LLM APIs are stochastic. While this would seem to doom any attempt at obtaining full outputs from the LLM, it is possible to counter certain types of stochasticity. In particular, we model stochastic API behavior as a collection of $n$ outputs $\boldsymbol{p}^{1},\boldsymbol{p}^{2},\ldots,\boldsymbol{p}^{n}$ from which the API randomly returns from. This might be the result of multiple instances of the LLM being run on different hardware which results in slightly different outputs. Whichever instance the API returns from determines which of the $n$ outputs we get. In order to determine which of the outputs the API returned from, let $p_{v-1}^{i}$ be the second highest token probability for output $\boldsymbol{p}^{i}$, and observe that $\log p^{i}_{v}-\log p^{i}_{v-1}=\log p^{i\prime}_{v}-\log p^{i\prime}_{v-1}$ for all outputs $\boldsymbol{p}^{i}$ and biased outputs $\boldsymbol{p}^{i\prime}$ where tokens $v$ and v1v-1 are not biased (proof in §A.3). Therefore, by biasing only k2k-2 tokens for each call, the API will return $p^{i\prime}_{v}$ and $p^{i\prime}_{v-1}$, which we can use to find $\log p^{i}_{v}-\log p^{i}_{v-1}$, which serves as a unique identifier for the distribution. Thus, after an average of nv/(k2)nv/(k-2) calls to the API we can collect the full set of probabilities for at least one of the outputs.
4 Fast, full outputs using the LLM image
The dominating factor in runtime for the algorithms in §3 is the vocabulary size $v$, which can be quite large (e.g., over 100 000100\,000 for OpenAI LLMs). We now introduce a preprocessing step that takes advantage of the low-dimensional LLM output space to obtain $O(d)$ versions of all the above algorithms. Since $d\ll v$ for many modern language models, this modification can result in multiple orders of magnitude speedups, depending on the LLM. For instance, the speedup for a model like pythia-70m (Biderman et al., 2023) would be $100\times$. The key to this algorithm is the observation from §2 that $d$ linearly independent outputs from the API constitute a basis for the whole output space (since the LLM’s image has dimension $d$). We can therefore collect these outputs $\boldsymbol{P}=\begin{bmatrix}\boldsymbol{p}^{1}&\boldsymbol{p}^{2}&\cdots&\boldsymbol{p}^{d}\end{bmatrix}\in\Delta_{v}^{d}$ as a preprocessing step in $O(vd)$ API calls using any of the above algorithms and $d$ unique prompts, and then use these to reconstruct the full LLM output after only $O(d)$ queries for each subsequent output.
To get a new full output $\boldsymbol{p}$, use any of the above algorithms to obtain $p_{1},p_{2},\ldots,p_{d}$. Since $\boldsymbol{p}$ resides in a $d$-dimensional space spanned by the columns of $\boldsymbol{P}$, the rest of the values of $\boldsymbol{p}$ are fully determined by these first $d$ values. §B gives the details of how to solve for the remaining values. Thus we can retrieve $\boldsymbol{p}$ in only $O(d)$ API queries.
This $(v/d)\times$ speedup makes any method that relies on full model outputs significantly cheaper so long as the number of model outputs needed exceeds $d$. Such methods include model stealing (Tramèr et al., 2016; Krishna et al., 2019) which attempts to learn a model that exactly replicates the behavior of a target model, and LM inversion (Morris et al., 2023) which uses LLM outputs to reconstruct hidden prompts. Additionally, the preprocessing step can be computed once then shared between any number of clients, further diluting the cost of full outputs.
5 Discovering the embedding size of API-protected LLMs
Algorithm 2 Our algorithm for finding the model image and embedding size for an API-protected LLM. To get the embedding size of the model without finding the image, extract only the first $t$ token probabilities at each step. With API caching this takes $O(vd/k)$ time, or $O(d^{2}/k)$ to only extract the embedding size.
function GetImage( $\mathsf{API}:\mathcal{V}^{*}\times\mathcal{P}(\mathcal{V}\times\mathbb{R})\to\mathcal{P}(\mathcal{V}\times\mathbb{R})$ )
  for $t=1,2,\ldots$ do
   $\boldsymbol{L}_{t}=\operatorname{clr}\left(\textsc{Extract}([t],\mathcal{V},\mathsf{API})\right)\in\mathbb{R}^{t\times\lvert\mathcal{V}\rvert} \triangleright$ Get full logits for $t$ contexts
   if $\operatorname{rank}(\boldsymbol{L}_{t})=t-100$ then $\triangleright$ When the output rank stops increasing
     return $\boldsymbol{L}_{t-100}$ $\triangleright$ Return the model image. Embedding size is t100t-100.
   end if
  end for
end function
Assuming only the generic output layer described in figure˜2, it is possible to infer the embedding size $d$ of an API-protected LLM from its outputs alone. Since the model outputs occupy a $d$-dimensional subspace of $\Delta_{v}$, collecting $d$ linearly independent outputs $\boldsymbol{p}^{1},\boldsymbol{p}^{2},\ldots,\boldsymbol{p}^{d}$ from the LLM forms a basis for the LLM’s image. In other words, subsequent model outputs will be a linear combination of the first $d$ outputs. We can therefore discover the value of $d$ by collecting outputs one at a time until the number of linearly independent outputs in the collection (i.e., the rank) stops increasing, which will occur after we have collected $d$ prompts. We find that it suffices to slightly over-collect, e.g., $d+1000$ outputs. Algorithm 2 formalizes this procedure.
To find the rank of model outputs, we use the fact that a matrix with $d$ linearly independent columns will have $d$ non-zero singular values. We use singular value decomposition to obtain the singular values of the matrix $\boldsymbol{L}=\begin{bmatrix}\operatorname{clr}(\boldsymbol{p}^{1})&\operatorname{clr}(\boldsymbol{p}^{2})&\cdots&\operatorname{clr}(\boldsymbol{p}^{d+1000})\end{bmatrix}$ and observe the index at which the magnitude of the singular values drops to zero, which occurs at index $d$. In practice, numerical imprecision causes the magnitudes drop almost to zero.
Figure 4: The singular values of outputs from LLMs with various known and unknown embedding sizes $d$. For each model with known embedding size, there is a clear drop in magnitude at singular value index $d$, indicating the embedding size of the model. Using this observation, we can guess the embedding size of gpt-3.5-turbo to be 4096.
To validate our method, we collect next-token distributions (conditioned on unique, 1-token prompts) from several open-source LLMs from the Pythia family (Biderman et al., 2023) with embedding sizes ranging from 512 to 1024. For all these models, the singular values of the resulting output matrix drop precisely at index $d$, as shown in figure˜4.
To demonstrate our method’s effectiveness against API-protected LLMs, we use our algorithm from §3.3 to collect nearly 6,000 next-token distribution outputs from gpt-3.5-turbo [^4] We specifically use model version 0125, accessed February 1–19, 2024., a popular API-protected LLM whose embedding size is not publicly disclosed. We find that this model’s singular values drop between index 4,600 and 4,650, indicating that the embedding size of this model is at most this size. This predicted embedding size is somewhat abnormal, since LLM embedding sizes are traditionally set to powers of two. If this were the case for gpt-3.5-turbo, it would be reasonable to guess that the embedding size is $2^{12}=$4096$\$. We predict that our raw estimate of 4600–4650 is an overestimate of the true embedding size, since any abnormal outputs due to errors (whether in our own code or OpenAI’s) would only increase the dimensionality of the observed output space. For instance, if we inadvertently collected 504 corrupted outputs, then a model with embedding size 4096 would appear to have an embedding size of 4600.
Knowing the embedding size alone is insufficient to ascertain an LLM’s parameter count. However, since most known transformer-based LLMs with embedding size 4096 have around $7$ billion parameters, it is likely that gpt-3.5-turbo has a similar number of active parameters. Any other parameter count would result in either abnormally narrow or wide models, which often perform worse. The actual parameter count may be much higher if the model uses a “mixture-of-experts” architecture (Shazeer et al., 2017). Given the active development and decreasing cost of inference with gpt-3.5-turbo, it is possible that its size and architecture has changed over time. Fortunately, our method can be used to monitor these updates over time, alerting end-users when LLM providers change embedding size (and presumably therefore, model size).
6 Attributing model outputs and auditing model updates
An LLM image can serve as an identifying signature for the model, as shown in figure˜5, where the logit output from a pythia-70m checkpoint lies uniquely in the checkpoint’s image, and not in the image of the preceding or following checkpoints, nor the checkpoints of any other similar model. [^5] Open source models with detailed checkpoint information (e.g., Biderman et al., 2023; Liu et al., 2023) make this type of detailed analysis possible. This suggests that LLM images are highly sensitive to parameter changes, which makes sense because the intersection of two different $d$-dimensional spaces is an even lower-dimensional space. Thus, it is possible to determine precisely which LLM produced a particular full-vocabulary output using only API access to a set of LLMs, and notably, without knowing the exact inputs to the model.